### v1

# Robust Preference for Dynamical Dark Energy in DESI BAO and SN Measurements

Abstract
Recent Baryon Acoustic Oscillation (BAO) measurements released by DESI, when combined with Cosmic Microwave Background (CMB) data from Planck and two different samples of Type Ia supernovae (Pantheon-Plus and DESY5) reveal a preference for Dynamical Dark Energy (DDE) characterized by a present-day quintessence-like equation of state that crossed into the phantom regime in the past. A core ansatz for this result is assuming a linear Chevallier-Polarski-Linder (CPL) parameterization $w(a)=w_{0}+w_{a}(1-a)$ to describe the evolution of the DE equation of state (EoS). In this paper, we test if and to what extent this assumption impacts the results. To prevent broadening uncertainties in cosmological parameter inference and facilitate direct comparison with the baseline CPL case, we focus on 4 alternative well-known models that, just like CPL, consist of only two free parameters: the present-day DE EoS ($w_{0}$) and a parameter quantifying its dynamical evolution ($w_{a}$). We demonstrate that the preference for DDE remains robust regardless of the parameterization: $w_{0}$ consistently remains in the quintessence regime, while $w_{a}$ consistently indicates a preference for a dynamical evolution towards the phantom regime. This tendency is significantly strengthened by DESY5 SN measurements. By comparing the best-fit $\chi^{2}$ obtained within each DDE model, we notice that the linear CPL parameterization is not the best-fitting case. Among the models considered, the EoS proposed by Barboza and Alcaniz consistently leads to the most significant improvement.
I Introduction
One of the most undoubtedly fascinating and unforeseen discoveries of the past three decades is that the Universe is undergoing an accelerated phase of expansion. This was first argued in 1998 through observations of distant Type Ia Supernovae Riess et al. (1998); Perlmutter et al. (1999), and has since been corroborated by a wide variety of other probes Riess et al. (2001); Tegmark et al. (2004); Scranton et al. (2003); Tonry et al. (2003); Knop et al. (2003); Seljak et al. (2005); Feng et al. (2005); Riess et al. (2004); Astier et al. (2006); Eisenstein et al. (2005, 2007); Tegmark et al. (2006); Sahni and Starobinsky (2006); Wood-Vasey et al. (2007); Vikhlinin et al. (2009); Stern et al. (2010); Sherwin et al. (2011); Bennett et al. (2013); Hinshaw et al. (2013); Dawson et al. (2013); de Jong et al. (2013); Anderson et al. (2014); Weinberg et al. (2013); Beutler et al. (2014); Delubac et al. (2015); Betoule et al. (2014); Aubourg et al. (2015); Ross et al. (2015); Moresco et al. (2016a, b); Rubin and Hayden (2016); Alam et al. (2017); Abbott et al. (2016); Haridasu et al. (2017); Troxel et al. (2018); Scolnic et al. (2018); Aghanim et al. (2020a, b); Gómez-Valent (2019); Yang and Gong (2020); Choi et al. (2020); Aiola et al. (2020); Alam et al. (2021); Nadathur et al. (2020); Rose et al. (2020); Di Valentino et al. (2020a); Alam et al. (2021); Asgari et al. (2021); Tröster et al. (2021); Dutcher et al. (2021); Abbott et al. (2022); Moresco et al. (2022); Abbott et al. (2023a); Brout et al. (2022); Madhavacheril et al. (2024); Abbott et al. (2023b); Adame et al. (2024a); Lodha et al. (2024); Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024). [^1] For a few caveats, objections, and discussions surrounding this conclusion raised over the years, see Refs. Buchert (2000, 2001, 2008); Hunt and Sarkar (2010); Nielsen et al. (2016); Tutusaus et al. (2017); Dam et al. (2017); Colin et al. (2019); Desgrange et al. (2019); Koksbang (2019a, b); Heinesen (2023).
Since all known forces and components in nature would decelerate the expansion rate of the Universe, acceleration itself requires a physical mechanism beyond the Standard Model of fundamental interactions, able to counteract deceleration, inducing instead a Dark Energy (DE) phase where the dynamics is characterized by negative pressure with an effective Equation of State (EoS) $w<-1/3$ Sahni and Starobinsky (2000); Carroll (2001); Peebles and Ratra (2003); Padmanabhan (2003); Copeland et al. (2006); Caldwell and Kamionkowski (2009); Li et al. (2011); Martin (2012).
In the theoretical framework described by the standard $\Lambda$CDM model of cosmology, DE is parametrized by a positive cosmological constant term ($\Lambda$) in the Einstein equation with its energy density comprising the vast majority of the Universe’s energy budget. Despite its apparent simplicity, this interpretation is not free from conceptual problems and limitations Weinberg (1989); Krauss and Turner (1995); Sahni and Starobinsky (2000); Weinberg (2000); Padmanabhan (2003); Sahni (2002); Yokoyama (2003); Nobbenhuis (2006); Burgess (2015); Joyce et al. (2015); Bull et al. (2016); Wang et al. (2016a). Foremost, plugging a positive cosmological constant component into the gravitational field equations described by General Relativity (GR) implies living in an asymptotically de Sitter universe, which seems to contrast with several theories/models of quantum gravity proposing instead an asymptotically anti-de Sitter universe Brustein and Steinhardt (1993); Witten (2000); Kachru et al. (2003); Polchinski (2006); Danielsson and Van Riet (2018). Secondly, it seems quite natural to question what made us so lucky to live precisely in the cosmic epoch when such a constant component came to be not only relevant but even dominant compared to other contributions such as matter, altering the expansion history of the Universe so prominently as to allow us to become easily aware of its existence and implications Zlatev et al. (1999); Pavon and Zimdahl (2005); Martin (2012); Velten et al. (2014). Finally, and most importantly, when it comes to the physical interpretations, anything that contributes to the energy density of the vacuum behaves akin to a cosmological constant, summing up within the energy-momentum tensor as $T_{\mu\nu}\propto g_{\mu\nu}\rho$ due to Lorentz invariance. Based on standard quantum field theory calculations, one would expect a zero-point energy density contribution of $\rho_{\rm vac}$, which, depending on the ultraviolet wavelength cutoff scale, is found to be somewhere between $10^{50}$ to $10^{120}$ orders of magnitude larger than what is inferred by cosmological data Weinberg (1989). This leads to one of the biggest disagreements between theoretical predictions and observations, arguably requiring a level of fine-tuning that appears to be well beyond what any current theories can realistically explain Dolgov (1997); Weinberg (1989); Straumann (1999); Padmanabhan (2003); Sola (2013). As a result, from a theoretical perspective, the nature of DE remains one of the biggest puzzles in modern physics, sustaining significant research interest in the high-energy physics community. [^2] This diffuse interest is reflected in the wide range of models – both within and beyond the standard cosmological constant – that have been proposed over the years. These include, for example, new (ultra)light fields and modifications to gravity. With no claims to completeness see, e.g., Refs. Amendola (2000); Kamenshchik et al. (2001); Capozziello (2002); Bento et al. (2002); Mangano et al. (2003); Farrar and Peebles (2004); Khoury and Weltman (2004); Li (2004); Amendola et al. (2007); Hu and Sawicki (2007); Cognola et al. (2008); Nojiri et al. (2011); Zhang and Sasaki (2012); Rinaldi (2015a); Luongo and Quevedo (2014); Rinaldi (2015b); De Felice et al. (2016); Wang et al. (2016a); Josset et al. (2017); Burrage and Sakstein (2016); Sebastiani et al. (2017); Nojiri et al. (2017); Burrage and Sakstein (2018); Capozziello et al. (2018); Benisty and Guendelman (2018); Casalino et al. (2018); Yang et al. (2018a); Saridakis et al. (2018); Visinelli and Vagnozzi (2019); Langlois (2019); Benisty et al. (2019); Boshkayev et al. (2019); Heckman et al. (2019); D’Agostino (2019); Mukhopadhyay et al. (2020a); Mukhopadhyay and Majumdar (2019); Mukhopadhyay et al. (2020b); Vagnozzi et al. (2020); Akarsu et al. (2020); Saridakis (2020); Ruchika et al. (2023); Odintsov et al. (2020a, b); Oikonomou (2021a, b); Vagnozzi et al. (2021a); Solanki et al. (2021); Saridakis (2021); Arora et al. (2022); Capozziello et al. (2022); Narawade et al. (2022); D’Agostino et al. (2022); Oikonomou and Giannakoudi (2022); Belfiglio et al. (2023); Luciano and Giné (2023); Kadam et al. (2023); Ong (2023); Bernui et al. (2023); Luciano (2023); Giani and Piattella (2023); Belfiglio et al. (2024); Frion et al. (2023); Adil et al. (2023); Halder et al. (2024).
[…]
Despite these remarkable achievements, it is no exaggeration to say that at the time of writing, observations are inconclusive about the physical nature of DE. While too far-away deviations from the canonical $\Lambda$CDM model appear severely constrained Gariazzo et al. (2024), several alternative theoretical frameworks and phenomenological avenues featuring new physics in the dark sector of the model remain at the very least plausible.
Trying to summarize an otherwise very articulated debate, we can adhere to Occam’s razor principle and start considering one of the simplest hypotheses beyond the cosmological constant. This involves assuming that DE can be modeled as a generic fluid with a constant EoS, $w_{0}$. By leaving $w_{0}$ as a free parameter in the theoretical model, cosmological observations can constrain deviations from the cosmological constant value ($w_{0}=-1$). [^4] See, e.g., Refs Bean and Melchiorri (2002); Hannestad and Mortsell (2002); Said et al. (2013); Shafer and Huterer (2014); Zhang (2016); Moresco et al. (2016b); Xu and Zhang (2016); Wang et al. (2016b); Vagnozzi et al. (2017); Zhang (2017); Feng et al. (2018); Scolnic et al. (2018); Wang (2018); Sprenger et al. (2019); Poulin et al. (2018); Roy Choudhury and Naskar (2019); Abbott et al. (2019); Aghanim et al. (2020b); Wang (2021); Roy Choudhury and Hannestad (2020); Alam et al. (2021); Chudaykin et al. (2021); D’Amico et al. (2021); Vagnozzi et al. (2021b); Yang et al. (2021b); Di Valentino (2021); Abbott et al. (2023a); Tröster et al. (2021); Brieden et al. (2022); Grillo et al. (2020); Cao et al. (2022); Zhang et al. (2021); Colgáin et al. (2021); Teng et al. (2021); Krishnan et al. (2021); Nunes and Vagnozzi (2021); Bernardo et al. (2022); Vagnozzi et al. (2022); Bargiacchi et al. (2022); Moresco et al. (2022); Semenaite et al. (2023); Carrilho et al. (2023); Wang (2022); Koussour et al. (2023); Bernardo et al. (2023); Narawade and Mishra (2023); Hou et al. (2023); Escamilla et al. (2024a); Kumar et al. (2023); Bhagat et al. (2023); Mussatayeva et al. (2023); Hogg (2024); Adame et al. (2024b) for recent and not-so-recent discussions and constraints on the DE EoS from a variety of astrophysical and cosmological probes. In this regard, focusing exclusively on the Planck CMB data one might speculate about a preference for a phantom-like DE component ($w_{0}<-1$Aghanim et al. (2020b); Escamilla et al. (2024a). [^5] In recent years, the possibility that the DE EoS can be phantom in nature has gained substantial interest, as in principle a shift of $\sim 20\%$ towards $w_{0}<-1$ could already be enough to address the well-known Hubble tension Verde et al. (2019); Di Valentino et al. (2021b); Perivolaropoulos and Skara (2022); Abdalla et al. (2022); Hu and Wang (2023) – see, e.g., Refs. Vagnozzi (2020); Giarè (2024) for an overview, as well as for caveats surrounding this possibility. However, as extensively documented in the literature, this preference is not confirmed by independent CMB experiments such as ACT and SPT Aiola et al. (2020); Dutcher et al. (2021); Balkenhol et al. (2023); Giarè (2023), and – most importantly – it lacks consistent support from observations of the local universe. When combining CMB, BAO and SN data altogether, no convincing deviation from $w_{0}=-1$ is seen, possibly lending weight to the cosmological constant interpretation Escamilla et al. (2024a).
However, one may argue that simplicity may not always be a prerogative of nature. Pushing this approach forward, we can relax the assumption of a constant EoS and consider models where $w(a)$ varies with the universe’s expansion – here and after known as Dynamical Dark Energy (DDE). This possibility, along with the various proposed physical realizations, has also undergone extensive testing; see, e.g., Refs. Mainini et al. (2003); Alam et al. (2004); Sola and Stefancic (2006); Antoniadis et al. (2007); Yang et al. (2021b); Zhao et al. (2012); Solà Peracaula et al. (2018); Sola et al. (2017); Zhao et al. (2017); Yang et al. (2017, 2019a, 2019b); Sola Peracaula et al. (2019); Pan et al. (2019); Colgáin et al. (2021); Escamilla-Rivera and Nájera (2022); Zhao et al. (2020). From an observational standpoint, CMB data alone have limited capacity to constrain DDE models due to minimal effects left at the epoch of the last scattering surface and the increased number of cosmological parameters Di Valentino et al. (2017, 2020d, 2022). [^6] This difficulty is often referred to as geometrical degeneracy. At its core, the problem is that different combinations of late-time cosmic parameters can be adjusted in such a way that the acoustic angular scale $\theta_{s}$ – determined by the ratio of the comoving sound horizon at recombination to the comoving distance to last scattering – remains constant if both quantities change proportionally. Consequently, measurements based solely on this scale cannot provide strong constraints on (dynamical) DE parameters by themselves, unless perturbation-level effects and late-time data are also incorporated to break this degeneracy. Even in simple parametrizations aimed at minimizing the number of free degrees of freedom, CMB experiments produce constraints that are typically too broad to provide informative results. Therefore, local universe observations acquire primary importance.
A significant turning point in the study of DDE models is marked by the very recent BAO released by the Dark Energy Spectroscopic Instrument (DESI) Adame et al. (2024a); Lodha et al. (2024) and – albeit to a lesser extent – by SN distance moduli measurements resulting from the five-year observations of the Dark Energy Survey (DESY5) Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024). Before these updated data, no significant preference had ever emerged in favor of DDE models, certainly not to the extent required to challenge the baseline cosmological constant interpretation Gariazzo et al. (2024). Excitingly, when DESI BAO observations are combined with Planck CMB data and SN distance moduli measurements (either from the Pantheon-plus catalog Scolnic et al. (2022); Brout et al. (2022) or from DESY5 Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024)), they produce strong indications for DDE. Specifically, within the linear Chevallier-Polarski-Linder (CPL) parameterization of the DE EoS $w(a)=w_{0}+w_{a}(1-a)$ Chevallier and Polarski (2001); Linder (2003) – where $w_{0}$ represents the present-day DE EoS and $w_{a}$ quantifies the dynamical evolution – we observe a preference for $w_{0}>-1$ and $w_{a}<0$ at a statistical level ranging between 2.5 and 3.9 $\sigma$, depending on the specific combination of data involved Adame et al. (2024b).
Unexpectedly, these results have heated up the recent debate, fueling a multitude of re-analyses and re-interpretations Cortês and Liddle (2024); Patel and Amendola (2024); Orchard and Cárdenas (2024); Liu et al. (2024); Chudaykin and Kunz (2024); Notari et al. (2024); Gialamas et al. (2024); Wang et al. (2024b); Lodha et al. (2024); Wang and Piao (2024); Giarè et al. (2024b); Carloni et al. (2024); Colgáin et al. (2024); Tada and Terada (2024); Yin (2024); Luongo and Muccino (2024); Park et al. (2024); Wang et al. (2024a); Shlivko and Steinhardt (2024); Ye et al. (2024); Li et al. (2024). As “extraordinary claims require extraordinary evidence”, we certainly advise and exercise caution. However, even taking in mind all the possible caveats and limitations surrounding the first DESI data release, it is undeniable that a high statistical preference for DDE holds intrinsic interest – if confirmed, this would represent the first concrete evidence of new physics beyond the standard model of cosmology.
Given the potential implications of this result, it is natural to question its robustness. Barring any possible systematic issues in these datasets, we examine if and to what extent assuming a linear parameterization for the DE EoS might impact the current findings. Although the CPL parameterization has been demonstrated to match the background evolution of distances arising from exact DE equations of motion to an accuracy of approximately 0.1% for viable cosmologies over a wide range of physics, including scalar fields, modified gravity, and phase transitions, other parameterizations proposed over the years (which may deviate from CPL at both $z\ll 1$ and $z\gtrsim 1$) remain allowed by current observations. Testing these alternative models against new data can certainly represent a useful exercise to shed light on the role played by the parameterization itself. To be fair, the process has already begun with several independent groups actively engaged in this activity Hernández-Almada et al. (2024); Pourojaghi et al. (2024); Ramadan et al. (2024); Berghaus et al. (2024); Qu et al. (2024b); Notari et al. (2024); Adolf et al. (2024).
Given the vast number of parameterizations proposed over the years and recently analyzed in relation to the DESI data, a few warnings are in order. First and foremost, alternative parameterizations often introduce extra parameters compared to CPL. This is both a blessing and a curse: on the one hand, accounting for more degrees of freedom allows more flexibility in $w(a)$. On the other hand, this typically implies relaxing the overall constraining power due to the combined effects of degeneracies and correlations among parameters. Secondly, the physical interpretations of the parameters involved may differ from the two employed in the CPL model. This further compounds the comparison of the results, making it difficult to derive general guidelines on the preference toward DDE.
To overcome these difficulties, in this article, we test different DDE models while allowing for a fair comparison of the results. We restrict our attention to five well-known parameterizations that satisfy the following criteria: (i) they consist of two parameters to describe the evolution of $w(a)$, $w_{0}$ and $w_{a}$; (ii) these two parameters retain the same physical meaning as in the CPL parameterization; (iii) for the same combinations of pairs $w_{0}$ $w_{a}$, the resulting shape of $w(a)$ deviates from CPL either either near the present epoch or in the past, depending on the specific case.
The paper is organized as follows. In Sec. II we sketch the theoretical set-up of the gravitational equations and propose the DE parametrizations we wish to study. In Sec. III we describe the observational data and the methodology to constrain the proposed DE parametrizations. In Sec. IV we present the constraints on the resulting DE scenarios. Finally, in Sec. V we draw our general conclusions.
II Dynamical Dark Energy Models
Considering a DDE component in cosmological models produces changes both in the background dynamics and in the dynamics of cosmological perturbations.
Focusing on flat models where the DE EoS $w(a)$ can be described by a continuous function of the scale factor, the first Friedman equation reads
$H^{2}(a)=\frac{8\pi G}{3}\bigg{[}$ $\rho_{\rm r,0}\,a^{-4}+\rho_{\rm m,0}\,a^{-3}+$
$+\rho_{\rm DE,0}\,a^{-3}\exp\left(-3\int_{a_{0}=1}^{a}\frac{w_{\rm DE}(a^{\prime})}{a^{\prime}}\,da^{\prime}\right)\bigg{]}.$ (1)
Here $H(a)=\dot{a}/a$ is the Hubble parameter, the dot denotes the derivative with respect to physical time $t$, the subscript 00 indicates quantities evaluate at present while $\rho_{\rm r}$, $\rho_{\rm m}$, and $\rho_{\rm DE}$ are the energy densities in radiation, matter, and DE, respectively.
When it comes to cosmological perturbations, diffeomorphism invariance of GR requires fixing a gauge. In the synchronous gauge, the line element reads Ma and Bertschinger (1995)
$ds^{2}=a^{2}(\tau)\left[-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\right],$ (2)
where $d\tau=dt/a(t)$ is the conformal time, $\delta_{ij}$ and $h_{ij}$ are the unperturbed and perturbed spatial part of the metric tensors. For a fluid component $i$, the equations governing the dimensionless density perturbations $\delta_{i}=\delta\rho_{i}/\rho_{i}$ and the divergence of the $i$-th fluid velocity $\theta_{i}=i\kappa^{j}v_{j}$ in the Fourier space are Ma and Bertschinger (1995):
$\delta^{\prime}_{i}$ $=$ $-(1+w_{i})\,\left(\theta_{i}+\frac{h^{\prime}}{2}\right)-3\mathcal{H}\left(\frac{\delta P_{i}}{\delta\rho_{i}}-w_{i}\right)\delta_{i}$ (3)
$-9\mathcal{H}^{2}\left(\frac{\delta P_{i}}{\delta\rho_{i}}-c^{2}_{a,i}\right)(1+w_{i})\frac{\theta_{i}}{{\kappa}^{2}},$
$\theta^{\prime}_{i}$ $=$ $-\mathcal{H}\left(1-3\frac{\delta P_{i}}{\delta\rho_{i}}\right)\theta_{i}+\frac{\delta P_{i}/\delta\rho_{i}}{1+w_{i}}\,{\kappa}^{2}\,\delta_{i}-{\kappa}^{2}\sigma_{i},$ (4)
where the primes denote the derivative with respect to conformal time $\tau$, $h$ is the usual synchronous gauge metric perturbation, $\mathcal{H}(a)=a^{\prime}/a$ is the conformal Hubble parameter, $\kappa$ is the wavenumber in Fourier space, $\sigma_{i}$ stands for the anisotropic stress of the $i$-th fluid, and $c^{2}_{a,i}$ represents the adiabatic sound speed of the $i$-th fluid defined as $c_{a,i}^{2}=\dot{P}_{i}/\dot{\rho}_{i}$. In this work, we fix the squared sound speed of the DE component in the rest frame $c^{2}_{\rm s,DE}=\frac{\delta P_{\rm DE}}{\delta\rho_{\rm DE}}=1$ as broadly expected for the simplest DE models based on a single light minimally coupled scalar field with a canonical kinetic term.
Having outlined the background and perturbation dynamics, we now list and describe the five different two-parameter models employed for $w(a)$.
II.1 Chevallier-Polarski-Linder parametrization
The model proposed by Chevallier, Polarski and Linder Chevallier and Polarski (2001); Linder (2003) (CPL hereafter) can be regarded as the baseline parameterization used in most analyses focusing on DDE, including this work. In this scenario, the DE EoS reads
$w_{\rm DE}(a)=w_{0}+w_{a}\times\left(1-a\right).$ (5)
As already mentioned in the introduction, its advantages include a manageable 2-dimensional phase space, reduction to linear redshift behavior at low redshift, bounded behavior at high redshift, high accuracy in reconstructing various scalar field equations of state and the resulting distance-redshift relations up to 0.1% accuracy, good sensitivity to observational data, and a straightforward physical interpretation. The latter arises from its representation as the Taylor expansion of $w(a)$ around the present epoch $a\simeq a_{0}\equiv 1$ up to the first order: $w_{0}=w(a_{0})$ and $w_{a}=-\frac{dw(a)}{da}\bigg{|}_{a=a_{0}}$ which is the coefficient for the dynamical term.
II.2 Exponential parametrization
As a next step, we consider the exponential form for the DE EoS Dimakis et al. (2016); Pan et al. (2020b)
$w_{\rm DE}(a)=(w_{0}-w_{a})+w_{a}\times\exp\left(1-a\right).$ (6)
Up to the first order of the Taylor expansion, this description reduces to the CPL parameterization around $a\simeq a_{0}\equiv 1$. However, as $a$ moves far away from 1, the exponential form can introduce (small) deviations from the linear CPL regime without increasing the dimensionality of the parameter space Najafi et al. (2024).
II.3 Jassal-Bagla-Padmanabhan parametrization
The third parameterization studied in this work is the model proposed by Jassal-Bagla-Padmanabhan in Ref. Jassal et al. (2005) (JBP hereafter). In this case the DE EoS is
$w_{\rm DE}(a)=w_{0}+w_{a}\times a\left(1-a\right).$ (7)
It is characterized by the sum of a linear and a quadratic term in the scale factor. When $a^{2}$ is close to 1, the term $-w_{a}a^{2}$ becomes comparable to $w_{a}a$, thereby leading to expected differences at low redshift compared to CPL.
II.4 Logarithmic parametrization
We consider the following logarithmic form for the EoS:
$w_{\rm DE}(a)=w_{0}-w_{a}\times\ln a.$ (8)
To the best of our knowledge, this parameterization was originally introduced by G. Efstathiou in Ref. Efstathiou (1999) to capture the behaviour of a wide class of potential scalar field models of DE at low redshift $z\lesssim 4$. Here, with a fair amount of courage and following thorough stability tests, we extend this parameterization all the way up to $z\to\infty$. In principle, for some combinations of parameters, the logarithmic term can actually grow in absolute value and cause instabilities. [^7] To address this problem, in Ref. Yang et al. (2021b), some of us adopted flat priors $w_{a}\in[-3,0]$, thus removing positive values of $w_{a}$ a priori. In this work, we have performed additional stability tests, which revealed that numerical instabilities arise only when CMB data are considered on their own (we do not report these results for any of the models under study, as they are not informative). Without late-time data, DE parameters remain essentially unconstrained, and $w_{a}$ can acquire large positive values. Despite the logarithmic nature of the equation of state, for these values, the DE energy density does not remain negligible in the early Universe, triggering warnings and errors in the Boltzmann solver code CAMB Lewis et al. (2000); Howlett et al. (2012). In contrast, when late-time data are included, the DE parameters are significantly constrained in the region $w_{a}<0$, and no large positive values are allowed. As a result, in the allowed region of parameter space, the DE energy density remains negligible at early times and does not affect early Universe cosmology. However, given the current data constraints and the slow logarithmic growth, this is not the case. We find that the parameterization can be safely extended to high redshift because the DE contribution remains largely negligible compared to other components in the Universe’s energy budget.
II.5 Barboza-Alcaniz parametrization
The last (but as we shall see, not least) model involved in our analysis is the one proposed by Barboza and Alcaniz in Ref. Barboza and Alcaniz (2008) (referred to as BA hereafter). In this case, the dark energy equation of state (EoS) is characterized by the following functional form:
$w_{\rm DE}(a)=w_{0}+w_{a}\times\frac{1-a}{a^{2}+(1-a)^{2}}.$ (9)
This parameterization shows a linear behavior at low redshifts and remains well-behaved as $z\to\infty$, while allowing for deviations from the baseline CPL scenario.
Parameter Prior
$\Omega_{\mathrm{b}}h^{2}$ $[0.005,0.1]$
$\Omega_{\mathrm{c}}h^{2}$ $[0.01,0.99]$
$\log(10^{10}A_{\mathrm{s}})$ $[1.61,3.91]$
$n_{\mathrm{s}}$ $[0.8,1.2]$
$\tau$ $[0.01,0.8]$
$100\theta_{\mathrm{MC}}$ $[0.5,10]$
$w_{0}$ $[-3,1]$
$w_{a}$ $[-3,2]$
Table 1: Ranges for the flat prior distributions imposed on the free cosmological parameters in the analysis.
III Methods
In this section, we describe the statistical methodologies and observational datasets employed in our analysis.
III.1 Statistical Analyses
The cosmology resulting from all the five DDE models listed in Sec. II can be characterized by 7 free parameters: the physical baryon energy density $\Omega_{\mathrm{b}}h^{2}$, the physical cold dark matter energy density $\Omega_{\mathrm{c}}h^{2}$, the amplitude of the primordial scalar spectrum $A_{\mathrm{s}}$, its spectral index $n_{s}$, the optical depth to reionization $\tau$, the angular size of the sound horizon $\theta_{\rm{MC}}$, and the two free parameters describing the DE sector — i.e., the present-day value of the DE EoS $w_{0}$ and the parameter describing its dynamical evolution $w_{a}$. To compare the theoretical predictions against observations, we implement these models in five different modified versions of the publicly available cosmological code CAMB Lewis et al. (2000); Howlett et al. (2012) and explore the posterior distributions of the 7-dimensional parameter space by performing Markov Chain Monte Carlo (MCMC) analyses via the publicly available sampler Cobaya Lewis and Bridle (2002); Lewis (2013). We employ the fast dragging speed hierarchy implementation Neal (2005), which includes a convergence diagnostic following the Gelman-Rubin statistic Gelman and Rubin (1992). We adopt a threshold $R-1\lesssim 0.01$ for assessing chain convergence. In Tab. 1, we present the flat prior ranges on which the parameters are left to freely vary.
III.2 Datasets
The datasets involved in our analyses are:
  • Planck: Measurements of the Planck CMB temperature anisotropy and polarization power spectra, their cross-spectra, and the combination of the ACT and Planck lensing power spectrum. All CMB likelihoods employed in this work are listed below:
    • (i)
      Measurements of the power spectra of temperature and polarization anisotropies, $C_{\ell}^{TT}$, $C_{\ell}^{TE}$, and $C_{\ell}^{EE}$, at small scales ($\ell>30$), obtained by the Planck plik likelihood Aghanim et al. (2020b, d);
    • (ii)
      Measurements of the spectrum of temperature anisotropies, $C_{\ell}^{TT}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck Commander likelihood Aghanim et al. (2020b, d);
    • (iii)
      Measurements of the spectrum of E-mode polarization, $C_{\ell}^{EE}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck SimAll likelihood Aghanim et al. (2020b, d);
    • (iv)
      Reconstruction of the spectrum of the lensing potential, obtained by the Planck PR4 NPIPE data release Carron et al. (2022) used in combination with ACT-DR6 lensing likelihood Madhavacheril et al. (2024); Qu et al. (2024a). [^8] The NPIPE lensing map Carron et al. (2022) covers CMB angular scales in the range $100\leq\ell\leq 2048$ using the quadratic estimator and re-processing Planck time-ordered data with several improvements, including around 8% more data compared to the plik-lensing likelihood. Notice also that NPIPE and ACT-DR6 measurements explore distinct angular scales and overlap only across a portion of the sky. Additionally they have different noise levels and instrument-related systematics. Therefore they can be regarded as nearly independent lensing measurements.
  • DESI: baryon acoustic oscillations (BAO) measurements extracted by observations of galaxies & quasars Adame et al. (2024a), and Lyman-$\alpha$ Adame et al. (2024c) tracers from the first year of observations using the Dark Energy Spectroscopic Instrument (DESI). These include measurements of the transverse comoving distance, the Hubble horizon, and the angle-averaged distance as summarized in Tab. I of Ref. Adame et al. (2024b).
  • PantheonPlus: Distance moduli measurements of 1701 light curves of 1550 spectroscopically confirmed type Ia SN sourced from eighteen different surveys, gathered from the Pantheon-plus sample Scolnic et al. (2022); Brout et al. (2022).
  • DESY5: Distance moduli measurements of 1635 Type Ia SN covering the redshift range of $1.130.10<z<$ 1.13 that have been collected during the full five years of the Dark Energy Survey (DES) Supernova Program Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024), along with 194 low-redshift SN in the redshift range of $0.10.025<z<$ 0.1 which are in common with the Pantheon-plus sample Scolnic et al. (2022); Brout et al. (2022).
IV Results
In this section, we present the observational constraints on the five DDE models considered in this article. We discuss the results model by model, testing each case against three different data combinations: Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5. We make no secret that due to the large number of analyzed models and the similarity of the results obtained, the following discussion may appear somewhat repetitive (though necessary). Therefore, readers interested in the results of specific models can find the numerical constraints, two-dimensional correlations, and one-dimensional posterior distribution functions of key parameters as follows:
  • Table 2 and Figure 1 summarize the numerical constraints and parameter correlations for the baseline CPL case (5). The results for this case are detailed in Sec. IV.1.
  • Table 3 and Figure 2 present the results for the exponential parameterization (6), discussed in Sec. IV.2.
  • Table 4 and Figure 3 provide the results for the JBP EoS (7), discussed in Sec. IV.3.
  • Table 5 and Figure 4 summarize the results for the logarithmic parameterization (8), detailed in Sec. IV.4.
  • Table 6 and Figure 5 present the results for the BA parameterization (9), discussed in Sec. IV.5.
Conversely, readers interested in a comprehensive overview of the results, their interpretation, and implications can refer directly to Sec. V.
IV.1 Results for the CPL parameterization
The numerical constraints obtained by adopting a baseline CPL parametrization are given in Tab. 2. Fig. 1 displays key parameters that characterize this model (i.e., the present-day value of the EoS $w_{0}$ and the parameter quantifying its redshift evolution $w_{a}$) as well as their correlations with other cosmological parameters of intrinsic interest for the late-time expansion history of the Universe such as the Hubble constant $H_{0}$, the present-day matter fractional energy density $\Omega_{m}$, and the matter clustering parameter $S_{8}$.
Figure 1: CPL parametrization (5) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
We recover all the results discussed in the recent DESI release paper Adame et al. (2024b). For Planck+DESI, the constraints favor a present-day quintessence EoS with $w_{0}=-0.44^{+0.34}_{-0.21}$ at 68% CL [^9] Hereafter, constraints will always be quoted at 68% CL unless otherwise specified., showing a notable shift away from $w_{0}=-1$. On the other hand, $w_{a}=-1.81^{+0.37}_{-1.1}$ provides hints of dynamical evolution towards the phantom regime.
The addition of PantheonPlus significantly refines the constraints on the parameter space, reducing the error bars on the DE parameters by up to a factor of 5. Although $w_{0}$ shifts towards $-1$, it remains strictly in the quintessence regime: $w_{0}=-0.876\pm 0.045$. Consistent with DESI 2024 findings Lodha et al. (2024), the mean value of $w_{a}$ increases compared to Planck+DESI, now reading $w_{a}=-0.51^{+0.20}_{-0.17}$. This boosts the evidence for a past phantom-like DDE component to above 95% CL, see also Fig. 1.
When replacing PantheonPlus with DESY5 type Ia SN, we observe a shift of $w_{0}$ away from $-1$, resulting in $w_{0}=-0.726\pm 0.069$. This places $w_{0}$ deep in the quintessence regime, see also Fig. 1. Similarly, $w_{a}=-1.05^{+0.34}_{-0.28}$ is found to be non-zero at high statistical significance. Thus, the evidence for DDE remains stronger in the Planck+DESI+DESY5 case compared to Planck+DESI+PantheonPlus, and the cosmological constant case falls well outside the joint probability contours in the $w_{0}$-$w_{a}$ plane, as seen in Fig. 1.
Parameter Planck+DESI Planck+DESI+PantheonPlus Planck+DESI+DESY5
$\Omega_{\mathrm{c}}h^{2}$ $0.11993\pm 0.00098$ $0.11962\pm 0.00098$ $0.11978\pm 0.00098$
$\Omega_{\mathrm{b}}h^{2}$ $0.02238\pm 0.00014$ $0.02241\pm 0.00014$ $0.02239\pm 0.00014$
$100\theta_{\mathrm{MC}}$ $1.04092\pm 0.00029$ $1.04097\pm 0.00030$ $1.04094\pm 0.00030$
$\tau$ $0.0530\pm 0.0073$ $0.0545\pm 0.0071$ $0.0538\pm 0.0074$
$n_{\mathrm{s}}$ $0.9655\pm 0.0038$ $0.9662\pm 0.0038$ $0.3161\pm 0.0066$
$\log(10^{10}A_{\mathrm{s}})$ $3.040\pm 0.013$ $3.044\pm 0.013$ $3.042\pm 0.013$
$w_{0}$ $-0.44^{+0.34}_{-0.21}$ $-0.876\pm 0.045$ $-0.726\pm 0.069$
$w_{a}$ $-1.81^{+0.37}_{-1.1}$ $-0.51^{+0.20}_{-0.17}$ $-1.05^{+0.34}_{-0.28}$
$\Omega_{\mathrm{m}}$ $0.344^{+0.033}_{-0.027}$ $0.3088\pm 0.0067$ $0.3161\pm 0.0066$
$\sigma_{8}$ $0.791^{+0.021}_{-0.028}$ $0.8192\pm 0.0098$ $0.8130\pm 0.0091$
$S_{8}$ $0.846^{+0.016}_{-0.013}$ $0.8310\pm 0.0095$ $0.8345\pm 0.0097$
$H_{0}\mathrm{\,[km/s/Mpc]}$ $64.6^{+2.2}_{-3.3}$ $67.98\pm 0.72$ $67.22\pm 0.66$
$r_{\mathrm{drag}}\mathrm{\,[Mpc]}$ $147.11\pm 0.24$ $147.16\pm 0.23$ $147.14\pm 0.23$
$\Delta\chi^{2}$ $-6.8$ $-8.4$ $-15.2$
Table 2: CPL Parametrization (5) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{CPL}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.2 Results for the Exponential parametrization
We present the constraints obtained by assuming an exponential parametrization for the DE EoS in Tab. 3. In Fig. 2, we show the one-dimensional posterior distribution functions and the two-dimensional marginalized contours for the key cosmic parameters.
As usual, we test the model against three different combinations of data involving the DESI BAO measurements. Focusing on the minimal Planck+DESI combination, we find $w_{0}=-0.50\pm 0.27$ – significantly different from $-1$ and deep in the quintessence regime. Similarly, $w_{a}=-1.40^{+0.75}_{-0.62}$ is almost $2\sigma$ away from the non-dynamical $w_{a}=0$ case, lending weight to the Planck+DESI preference for DDE.
[…]
When we focus on the Planck+DESI+DESY5 data combination, the evidence for DDE becomes significantly more pronounced. $w_{0}=-0.804^{+0.045}_{-0.051}$ remains strictly in the quintessence regime, while $w_{a}=-0.71^{+0.23}_{-0.19}$ is $3\sigma$ away from the non-dynamical $w_{a}=0$ scenario; see again Fig. 2.
Figure 2: Exponential parameterization (6) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
Overall, in terms of constraints on cosmic parameters, these results are in agreement with those derived for the CPL parametrization in the previous section, underscoring the resilience of the evidence for DDE and relieving concerns about dependence on the model for these particular results.
Parameter Planck+DESI Planck+DESI+PantheonPlus Planck+DESI+DESY5
$\Omega_{\mathrm{c}}h^{2}$ $0.1201\pm 0.0010$ $0.11962\pm 0.00099$ $0.1198\pm 0.0010$
$\Omega_{\mathrm{b}}h^{2}$ $0.02237\pm 0.00014$ $0.02241\pm 0.00014$ $0.02239\pm 0.00014$
$100\theta_{\mathrm{MC}}$ $1.04089\pm 0.00030$ $1.04097\pm 0.00030$ $1.04094\pm 0.00030$
$\tau$ $0.0525\pm 0.0073$ $0.0547\pm 0.0073$ $0.0539\pm 0.0074$
$n_{\mathrm{s}}$ $0.9650\pm 0.0038$ $0.9663\pm 0.0038$ $0.9657\pm 0.0039$
$\log(10^{10}A_{\mathrm{s}})$ $3.040\pm 0.013$ $3.044\pm 0.013$ $3.042\pm 0.013$
$w_{0}$ $-0.50\pm 0.27$ $-0.820\pm 0.064$ $-0.804^{+0.045}_{-0.051}$
$w_{a}$ $-1.40^{+0.75}_{-0.62}$ $-0.77\pm 0.28$ $-0.71^{+0.23}_{-0.19}$
$\Omega_{\mathrm{m}}$ $0.352\pm 0.035$ $0.3088\pm 0.0070$ $0.3157\pm 0.0066$
$\sigma_{8}$ $0.788^{+0.026}_{-0.029}$ $0.8186\pm 0.0094$ $0.8150\pm 0.0092$
$S_{8}$ $0.852\pm 0.017$ $0.8304\pm 0.0097$ $0.8360\pm 0.0098$
$H_{0}\mathrm{\,[km/s/Mpc]}$ $64.0^{+2.9}_{-3.5}$ $67.98\pm 0.72$ $67.28\pm 0.65$
$r_{\mathrm{drag}}\mathrm{\,[Mpc]}$ $147.08\pm 0.24$ $147.16\pm 0.23$ $147.12\pm 0.24$
$\Delta\chi^{2}$ $-6.9$ $-7.8$ $-15.2$
Table 3: Exponential Parametrization (6) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{exp}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.3 Results for the JBP parametrization
The numerical constraints for the JBP parametrization are given in Tab. 4, while the marginalized probability contours for the usual parameters are shown in Fig. 3.
When considering Planck+DESI, unlike the other parametrizations described so far (e.g., CPL and exponential form), $w_{a}$ remains unbounded and an upper limit $w_{a}<0.648$ can be derived at 95% CL. Conversely, $w_{0}$ remains in the quintessence regime ($w_{0}=-0.79^{+0.31}_{-0.14}$), confirming the overall tendency for a present-day quintessence EoS.
When considering PantheonPlus in combination with Planck and DESI, we get $w_{a}=-1.50\pm 0.57$ – non-null at more than $2.6\sigma$. Additionally, the constraints on $w_{0}=-0.767\pm 0.086$ are narrowed down within the quintessence portion of the parameter space as seen in Fig. 3. Thus, effectively, evidence of DDE is confirmed for this parametrization as well.
Finally, we replace PantheonPlus with DESY5 SN measurements. In this case, we obtain $w_{0}=-0.641^{+0.095}_{-0.067}$ and $w_{a}=-2.12^{+0.38}_{-0.68}$, confirming that the evidence of DDE becomes much more pronounced with DESY5. This evidence reaches a statistical significance $\gtrsim 3\sigma$. Having said that, comparing Fig. 3 with the respective triangular plots of the other parameterizations, we notice that for this model, the uncertainties remain much broader, especially for the parameters describing DE EoS. This suggests that this parameterization might not be the ideal fitting model for current observations, hinting at some possible internal tension compared to the other cases. This concern is confirmed when comparing the difference between the best-fit $\chi^{2}$ obtained within each DDE model and the best-fit $\chi^{2}$ obtained within $\Lambda$CDM. Indeed, this model consistently leads to the smallest improvement over $\Lambda$CDM across all data and DDE models.
Figure 3: JBP parametrization (7) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
Parameter Planck+DESI Planck+DESI+PantheonPlus Planck+DESI+DESY5
$\Omega_{\mathrm{c}}h^{2}$ $0.11940\pm 0.00098$ $0.11934\pm 0.00096$ $0.11947\pm 0.00094$
$\Omega_{\mathrm{b}}h^{2}$ $0.02242\pm 0.00013$ $0.02243\pm 0.00014$ $0.02242\pm 0.00014$
$100\theta_{\mathrm{MC}}$ $1.04099\pm 0.00029$ $1.04102\pm 0.00029$ $1.04100\pm 0.00029$
$\tau$ $0.0558\pm 0.0075$ $0.0564\pm 0.0073$ $0.0557\pm 0.0072$
$n_{\mathrm{s}}$ $0.9668\pm 0.0038$ $0.9670\pm 0.0038$ $0.9667\pm 0.0038$
$\log(10^{10}A_{\mathrm{s}})$ $3.046\pm 0.014$ $3.048\pm 0.013$ $3.047\pm 0.013$
$w_{0}$ $-0.79^{+0.31}_{-0.14}$ $-0.767\pm 0.086$ $-0.641^{+0.095}_{-0.067}$
$w_{a}$ $<0.648$ $-1.50\pm 0.57$ $-2.12^{+0.38}_{-0.68}$
$\Omega_{\mathrm{m}}$ $0.304^{+0.023}_{-0.019}$ $0.3096\pm 0.0068$ $0.3180\pm 0.0065$
$\sigma_{8}$ $0.822^{+0.019}_{-0.025}$ $0.8151\pm 0.0093$ $0.8083\pm 0.0086$
$S_{8}$ $0.826^{+0.012}_{-0.011}$ $0.8279\pm 0.0093$ $0.8321\pm 0.0093$
$H_{0}\mathrm{\,[km/s/Mpc]}$ $68.6^{+1.9}_{-2.8}$ $67.83\pm 0.71$ $66.96\pm 0.64$
$r_{\mathrm{drag}}\mathrm{\,[Mpc]}$ $147.20\pm 0.23$ $147.21\pm 0.23$ $147.19\pm 0.22$
$\Delta\chi^{2}$ $-5.6$ $-6.4$ $-16.0$
Table 4: JBP parametrization (7) – 68% CL constraints and 95% CL upper limits on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{JBP}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.4 Results for the Logarithmic parametrization
Tab. 5 summarizes the constraints on the model where the DE EoS is described by the logarithmic parametrization. Fig. 4 displays the usual marginalized contours on relevant parameters.
Starting with Planck+DESI, $w_{0}=-0.48^{+0.28}_{-0.33}$ is confined to the quintessence regime at more than 68% CL, while $w_{a}=1.33^{+0.79}_{-0.56}$ is constrained to be different from zero at more than $2.3\sigma$ – confirming once more the preference for DDE in Planck+DESI.
When PantheonPlus is added to Planck+DESI, we find $w_{0}=-0.843\pm 0.055$; i.e., shifted towards $w_{0}=-1$ although with error bars smaller by a factor of 5. However, also in this parameterization, $w_{0}$ is preferred to be in the quintessence regime, excluding $w_{0}=-1$ at more than $2.8\sigma$. Similarly, the result on $w_{a}=-0.53^{+0.22}_{-0.18}$ confirms the overall preference for DDE, see Fig. 4.
Considering DESY5 in place of PantheonPlus, the constraints on $w_{0}$ and $w_{a}$ change to $w_{0}=-0.763^{+0.054}_{-0.062}$ and $w_{a}=-0.72^{+0.25}_{-0.19}$, respectively. As a result, $w_{0}$ remains in the quintessence regime at more than 95% CL, while $w_{a}$ is found to be non-zero at almost $3\sigma$. It is noteworthy that the evidence of DDE is consistently more pronounced in the presence of DESY5 compared to PantheonPlus, see Fig. 4.
Figure 4: Logarithmic parametrization (8) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
Parameter Planck+DESI Planck+DESI+PantheonPlus Planck+DESI+DESY5
$\Omega_{\mathrm{c}}h^{2}$ $0.1201\pm 0.0010$ $0.11964\pm 0.00099$ $0.11986\pm 0.00099$
$\Omega_{\mathrm{b}}h^{2}$ $0.02237\pm 0.00014$ $0.02240\pm 0.00014$ $0.02239\pm 0.00014$
$100\theta_{\mathrm{MC}}$ $1.04090\pm 0.00030$ $1.04095\pm 0.00029$ $1.04094\pm 0.00029$
$\tau$ $0.0520\pm 0.0073$ $0.0542\pm 0.0073$ $0.0535\pm 0.0073$
$n_{\mathrm{s}}$ $0.9648\pm 0.0039$ $0.9661\pm 0.0038$ $0.9657\pm 0.0038$
$\log(10^{10}A_{\mathrm{s}})$ $3.039\pm 0.013$ $3.043\pm 0.013$ $3.042\pm 0.013$
$w_{0}$ $-0.48^{+0.28}_{-0.33}$ $-0.843\pm 0.055$ $-0.763^{+0.054}_{-0.062}$
$w_{a}$ $-1.33^{+0.79}_{-0.56}$ $-0.53^{+0.22}_{-0.18}$ $-0.72^{+0.25}_{-0.19}$
$\Omega_{\mathrm{m}}$ $0.346^{+0.030}_{-0.035}$ $0.3086\pm 0.0067$ $0.3156\pm 0.0066$
$\sigma_{8}$ $0.792\pm 0.025$ $0.8189\pm 0.0097$ $0.8142\pm 0.0092$
$S_{8}$ $0.848\pm 0.016$ $0.8305\pm 0.0095$ $0.8350\pm 0.0094$
$H_{0}\mathrm{\,[km/s/Mpc]}$ $64.6^{+2.8}_{-3.1}$ $68.01\pm 0.71$ $67.29\pm 0.66$
$r_{\mathrm{drag}}\mathrm{\,[Mpc]}$ $147.07\pm 0.23$ $147.17\pm 0.24$ $147.12\pm 0.23$
$\Delta\chi^{2}$ $-6.5$ $-9.3$ $-14.8$
Table 5: Logarithmic parametrization (8) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{log}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.5 Results for the BA parametrization
The observational constraints for the last model analyzed in this work – the BA parametrization – are given in Tab. 6. As usual, we illustrate the correlations among the key cosmic parameters in Fig. 5.
Combining Planck with DESI, we get $w_{0}=-0.39^{+0.30}_{-0.34}$, approaching $-1/3$ and approximately $2\sigma$ away from $w_{0}=-1$. Additionally, $w_{a}=-1.07^{+0.55}_{-0.43}$ is significantly different from $w_{a}=0$, confirming the preference for DDE in a similar fashion to other parameterizations discussed throughout the manuscript.
The inclusion of PantheonPlus gives $w_{0}=-0.848\pm 0.054$ (deep in the quintessence regime) and $w_{a}=-0.38^{+0.15}_{-0.13}$ (non-zero at more than $2\sigma$). Thus, for Planck+DESI+PantheonPlus, evidence of dynamical dark energy is clearly indicated, consistent with all the other parameterizations described so far.
In the case of Planck+DESI+DESY5, $w_{0}=-0.770\pm 0.057$ shifts further away from $-1$, strengthening the preference for a quintessence EoS. Meanwhile, $w_{a}=-0.51^{+0.16}_{-0.14}$ is found to be non-zero at more than $3\sigma$, as visualized in Fig. 5.
Interestingly, when comparing the difference between the best-fit $\chi^{2}$ obtained within each DDE model and the best-fit $\chi^{2}$ obtained within $\Lambda$CDM, this model consistently leads to the most significant improvement over $\Lambda$CDM across all three data combinations analyzed. It performs better than the CPL parameterization, as seen by comparing the last line in Tab.2 and the last line in Tab.6.
Figure 5: BA parametrization (9) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
Parameter Planck+DESI Planck+DESI+PantheonPlus Planck+DESI+DESY5
$\Omega_{\mathrm{c}}h^{2}$ $0.1201\pm 0.0010$ $0.11963\pm 0.00099$ $0.1198\pm 0.0010$
$\Omega_{\mathrm{b}}h^{2}$ $0.02237\pm 0.00014$ $0.02240\pm 0.00014$ $0.02239\pm 0.00014$
$100\theta_{\mathrm{MC}}$ $1.04090\pm 0.00029$ $1.04097\pm 0.00030$ $1.04095\pm 0.00029$
$\tau$ $0.0523\pm 0.0073$ $0.0544\pm 0.0074$ $0.0539\pm 0.0074$
$n_{\mathrm{s}}$ $0.9649\pm 0.0039$ $0.9663\pm 0.0039$ $0.9658\pm 0.0038$
$\log(10^{10}A_{\mathrm{s}})$ $3.039\pm 0.013$ $3.044\pm 0.014$ $3.043\pm 0.013$
$w_{0}$ $-0.39^{+0.30}_{-0.34}$ $-0.848\pm 0.054$ $-0.770\pm 0.057$
$w_{a}$ $-1.07^{+0.55}_{-0.43}$ $-0.38^{+0.15}_{-0.13}$ $-0.51^{+0.16}_{-0.14}$
$\Omega_{\mathrm{m}}$ $0.357^{+0.033}_{-0.040}$ $0.3084\pm 0.0069$ $0.3155\pm 0.0066$
$\sigma_{8}$ $0.783\pm 0.028$ $0.8189\pm 0.0097$ $0.8138\pm 0.0093$
$S_{8}$ $0.852\pm 0.017$ $0.8302\pm 0.0095$ $0.8344\pm 0.0098$
$H_{0}\mathrm{\,[km/s/Mpc]}$ $63.6\pm 3.3$ $68.03\pm 0.73$ $67.30\pm 0.67$
$r_{\mathrm{drag}}\mathrm{\,[Mpc]}$ $147.07\pm 0.24$ $147.17\pm 0.23$ $147.13\pm 0.24$
$\Delta\chi^{2}$ $-8.7$ $-9.4$ $-16.2$
Table 6: BA parametrization (9) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{BA}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
V Discussions and Conclusions
[…]
In this paper, we tested whether and to what extent the preference for a present-day quintessence equation of state that evolves towards the phantom regime depends on the parameterization adopted to describe its dynamical behavior. To avoid broadening uncertainties in cosmological parameters and facilitate direct comparison with the baseline CPL case, we focused on some well-known alternative models: the exponential, Jassal-Bagla-Padmanabhan, logarithmic, and Barboza-Alcaniz parameterizations for the equation of state. Like the CPL model, all these parameterizations consist of only two free parameters: the present-day value of the equation of state ($w_{0}$) and a parameter quantifying its dynamical evolution ($w_{a}$). However, they allow for deviations from linear behavior at both high and low redshifts. Therefore, given the same pair of values ($w_{0}$, $w_{a}$), different late-time expansion histories are obtained within the four models, thereby affecting cosmological observables differently.
To assess whether the preference for a dynamical dark energy component characterized by $w_{0}>-1$ and $w_{a}<0$ remains a robust prediction of the data, we tested these models against the most recent high and low redshift observations: the Planck 2018 CMB measurements, DESI BAO, as well as PantheonPlus and DESY5 SN measurements. For all the dataset combinations explored – i.e., Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 – we find that $w_{0}$ consistently remains in the quintessence regime. Additionally, the constraints on $w_{a}$ consistently indicate a preference for a dynamical evolution that crossed into the phantom regime ($w_{a}<0$). Therefore, our findings confirm the DESI results, regardless of the parameterization adopted to describe the dynamics of the dark energy sector.
Figure 6: Summary Plot – two-dimensional marginalized contours in the ($w_{0}$, $w_{a}$) plane for all models and datasets analyzed in this study.
Notably, convincing hints of a dynamical evolution of the equation of state are found even with just Planck+DESI. As clearly seen in Fig. 6 – which summarizes the results for the different models – the pair $w_{0}=-1$ and $w_{a}=0$ (corresponding to the standard cosmological constant model of structure formation, $\Lambda$CDM) always falls outside the 95% confidence level contour.
However, the real step forward in terms of preference for dynamical dark energy comes when we consider Type Ia supernovae. Including distance moduli measurements gathered from the PantheonPlus catalog, the error bars on $w_{0}$ and $w_{a}$ tighten by a factor of 5 compared to Planck+DESI alone. The contours on $w_{0}$ significantly shrink within the quintessence portion of the parameter space $w_{0}>-1$, while the contours on $w_{a}$ significantly reduce within the $w_{a}<0$ region. Replacing PantheonPlus data with DESY5 SN measurements, the preference for dynamical dark energy becomes substantially more significant, to the point where it is not an exaggeration to refer to it as evidence rather than mere preference. This is again clearly illustrated in Fig. 6: for all models, the constraints shift further away from a cosmological constant, which always falls well outside the 95% marginalized probability contours.
At first glance, Fig. 6 also reveals that the contours in the $w_{0}$-$w_{a}$ plane show similar trends for all four parameterizations (including the baseline CPL case), especially when SN measurements are included in the analysis. This simultaneously underscores the intrinsic robustness of the preference for dynamical dark energy as reported by the DESI BAO and SN measurements and its resilience against different parameterizations. Given these results, there is solid ground to conclude that the choice of parameterization has a minimal impact.
Last but not least, we examined statistical metrics to quantify the extent to which the different parameterizations analyzed in this study are successful in explaining observations. Specifically, for each model and data combination, we report the difference between the best-fit $\chi^{2}$ obtained within each dynamical dark energy model and the best-fit $\chi^{2}$ obtained within $\Lambda$CDM.
Once more, all models exhibit similar trends: for Planck+DESI, we consistently observe an improvement in the fit over $\Lambda$CDM, with $\Delta\chi^{2}$ ranging from $-5.6$ to $-8.7$, depending on the specific model. This improvement in the fit is further enhanced when PantheonPlus SN measurements are included ($\Delta\chi^{2}$ ranges from $-6.4$ to $-9.4$) and is substantially increased – by up to a factor of $\sim 2$ – when adopting DESY5 SN data (in this case, $\Delta\chi^{2}$ ranges from $-14.8$ to $-16.2$). This trend follows the overall preference for dynamical dark energy discussed thus far. Interestingly, the linear CPL parameterization is never the best-fitting model. The equation of state proposed by Barboza-Alcaniz, given by Eq. (9), consistently leads to the most significant improvement in $\Delta\chi^{2}$ over $\Lambda$CDM across all three data combinations analyzed. Conversely, the Jassal-Bagla-Padmanabhan parameterization, given by Eq. (7), shows the smallest improvement in fit compared to $\Lambda$CDM among the models considered. These results suggest that current data are approaching a precision that could enhance our understanding of the physical nature behind the apparent dynamical behavior of dark energy, should future surveys and data releases confirm these findings.
Acknowledgements.
W.G. acknowledges support from the Lancaster–Sheffield Consortium for Fundamental Physics through the Science and Technology Facilities Council (STFC) grant ST/X000621/1. S.P. acknowledges the financial support from the Department of Science and Technology (DST), Govt. of India under the Scheme “Fund for Improvement of S&T Infrastructure (FIST)” (File No. SR/FST/MS-I/2019/41). E.D.V acknowledges support from the Royal Society through a Royal Society Dorothy Hodgkin Research Fellowship. This article is based upon work from the COST Action CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics” (CosmoVerse), supported by COST (European Cooperation in Science and Technology).